OpenStudy (anonymous):

Compound interest: How long does it take a 4000 dollar investment to double if it earns interest at the annual rate of 8%, compounded monthly?

OpenStudy (amistre64):

you will want to use your compound formula for this

OpenStudy (amistre64):

what are do you have written out as a formula so far?

OpenStudy (anonymous):

\[A=P(1+r/n)^nt\]

OpenStudy (anonymous):

Sorry that is incorrect I meant A=P[1+(r/n)]^nt

OpenStudy (amistre64):

good, and since we want to have A = 2P... we can sub it in divide of the P to start with, then log to reveal the exponent ... then divide off again

OpenStudy (amistre64):

2P= P(1+r/n)^nt 2=(1+r/n)^nt log(2)=nt log(1+r/n) considering this is just the form: a = bt, how do we solve for t?

OpenStudy (anonymous):

a=bt t=a/b

OpenStudy (amistre64):

yep a=log(2) b= n log(1+r/n) so we know all the parts, therefore we know t

OpenStudy (anonymous):

I do not understand

OpenStudy (amistre64):

you know that t = a/b ... and i just defined a and b ... what is not to understand?

OpenStudy (anonymous):

Can I just do this: 8000 = 4000[1+(0.08/12)]^12t

OpenStudy (amistre64):

that doesnt solve for t tho, but those are the proper placement for the information yes

OpenStudy (anonymous):

Then to solve for t i can write: a=log(2) b= n log(1+r/n) t = [log(2)]/[nlog(1+r/n)] ?

OpenStudy (amistre64):

yes, after working some algebra to get at 't', we can either work it in a general manner like that; or work the same process but with the specific values in it to start with ... either way. i prefer the general approach, otherwise we introduce errors when approximating the other values

OpenStudy (amistre64):

t= log(2)/(12 log(1+.08/12))

OpenStudy (amistre64):

so 8 months, plus part of another, so 9 months if we want to keep it in whole units

OpenStudy (amistre64):

years ... that is

OpenStudy (anonymous):

I got t = 8.6931

OpenStudy (amistre64):

8.69319 years... 8 years and .69319(12) months 8.32

OpenStudy (anonymous):

So about 8.32 years

OpenStudy (amistre64):

about 8.69 years more precisely .. 8 years and 8.32 months

OpenStudy (anonymous):

Ok makes sense. Thanks

OpenStudy (amistre64):

good luck :)